# Dataset Distillation via Knowledge Distillation: Towards Efficient Self-Supervised Pre-training of Deep Networks

### v1

@@ A.1 Additional Details for Experiments in Tables 1, 2, 3, 4, 5, 6, 7
Training the Teacher Model Using SSL We trained the teacher model using BarlowTwins [43] using the training setup ResNet18 specified in [10]. We used the Adam optimizer with batch size 256, learning rate 0.01, cosine annealing learning rate schedule, and weight decay $10^{-6}$. The feature dimension is 1024. Finally, we use the pre-projection head representation of the trained model for teacher representation, and its dimension is of size 512.
Training Expert Trajectories Using KD We trained 100 expert trajectories for each dataset with random initialization of the network for 20 epochs, using Stochastic Gradient Descent with learning rate 0.1, momentum 0.9, and weight decay 1e-4. Similar to other DD works [1; 21], we used depth 4 ConvNet for Tiny ImageNet and depth 3 ConvNet for both CIFAR 10 and CIFAR 100. We did not apply any augmentation except normalization, and did not apply the ZCA-whitening.
@@
Evaluation We use the models’ penultimate layer’s representations (as is standard in contrastive learning [43; 2]) of the downstream task’s training set and train a linear classifier, using LBFGS with 100 iterations and regularization weight $10^{-3}$. We then use the pre-projection head representations of the test set of the downstream task and evaluate using the aforementioned linear classifer to report the downstream test accuracy.
Using SimCLR for Obtaining Teacher Representation. We conducted an ablation study using a teacher model trained with SimCLR [2] instead of Barlow Twins [43] for CIFAR 10 and CIFAR 100. The experiment steps are similar to the ones in A.1. During the "Training the Teacher Model Using SSL" step , we used the Adam optimizer with batch size 512, learning rate 0.001, and weight decay $10^{-6}$ to train a ResNet18 along with a 2-layer linear projection head for 400 epochs. The projection head included Batch Normalization and ReLU after the first layer, and Batch Normalization after the second layer, projecting to 128 dimensions. Then, we used the pre-projection head representation of the trained model for getting the teacher representation of size 512. The other steps are the same as the one in A.1. Table 6 shows that MKDT consistently outperforms the random subset across all downstream datasets and various sizes of labeled data, highlighting the method’s generalizability to other contrastive learning methods.
@@ Appendix B Additional Comparison of Random Subset SSL Pre-Training to KRR-ST
This experiment is conducted in the setting of [21] for pre-training on CIFAR100. This experiment pre-trains on a distilled set / subset of 2% the size of CIFAR100 and then evaluates on downstream tasks by finetuning the entire network with the entire labeled dataset for the downstream task. KRR-ST compares with the random subset baseline by pretraining using SL. However, we use SSL on the random subset as a baseline instead. Here, we show that, in fact, even in the finetuning setting reported by [21], KRR-ST cannot outperform SSL pre-training on random real images. The baseline reported as ‘Random’ in [21] refers to SL pre-training as opposed to SSL pre-training.
@@
No-pretraining (from [21]) $64.95$±$\pm 0.21$ $87.34$±$\pm 0.13$ $34.66$±$\pm 0.39$ $19.43$±$\pm 0.14$ $18.46$±$\pm 0.11$ $22.31$±$\pm 0.22$ $58.75$±$\pm 0.41$
Random (Supervised Learning (from [21]) $65.23$±$\pm 0.12$ $87.55$±$\pm 0.19$ $33.99$±$\pm 0.45$ $19.77$±$\pm 0.21$ $18.18$±$\pm 0.21$ $21.69$±$\pm 0.18$ $59.31$±$\pm 0.27$
KRR-ST (from [21]) $66.81$±$\pm 0.11$ $88.72$±$\pm 0.11$ $41.54$±$\pm 0.37$ $28.68$±$\pm 0.32$ $25.30$±$\pm 0.37$ $26.39$±$\pm 0.08$ $67.88$±$\pm 0.18$
@@
From [38], we have that $L_{SSL}$ can be re-written as
@@ Conclusion
It is easy to see, from prior work in hard negative mining for contrastive SSL such as [30], that $P_{2}$ will have a different gradient across all batches from $P_{1}$ as $P_{2}$ only contrasts examples within the same class, whereas $P_{1}$ contrasts examples across different classes. Hence, $\theta_{SSL_{1}}\neq\theta_{SSL_{2}}$.
@@
[38] shows the $L_{SSL}$ can be rewritten as $L_{SSL}=-Tr(2M\theta\theta^{T})+Tr(M^{\prime}\theta^{T}\theta M^{\prime}\theta^{T}\theta)$
@@
From [38], we also have $\nabla_{\theta}(L_{SSL})=-4\theta M+4\theta M^{\prime}\theta^{T}\theta M^{\prime}$
@@
Appendix E Societal Impact

### v3

@@ 6 Conclusion
Acknowledgments This research was partially supported by the National Science Foundation CAREER Award 2146492, and an Amazon PhD fellowship.
@@ A.1 Additional Details for Experiments in Tables 1, 2, 3, 4, 5, 6, 7
Training the Teacher Model Using SSL We trained the teacher model using BarlowTwins (Zbontar et al., 2021) using the training setup ResNet18 specified in (Gedara Chaminda Bandara et al., 2023). We used the Adam optimizer with batch size 256, learning rate 0.01, cosine annealing learning rate schedule, and weight decay $10^{-6}$. The feature dimension is 1024. Finally, we use the pre-projection head representation of the trained model for teacher representation, and its dimension is of size 512.
Training Expert Trajectories Using KD We trained 100 expert trajectories for each dataset with random initialization of the network for 20 epochs, using Stochastic Gradient Descent with learning rate 0.1, momentum 0.9, and weight decay 1e-4. Similar to other DD works (Cazenavette et al., 2022; Lee et al., 2023), we used depth 4 ConvNet for Tiny ImageNet and depth 3 ConvNet for both CIFAR 10 and CIFAR 100. We did not apply any augmentation except normalization, and did not apply the ZCA-whitening.
@@
Evaluation We use the models’ penultimate layer’s representations (as is standard in contrastive learning (Zbontar et al., 2021; Chen et al., 2020)) of the downstream task’s training set and train a linear classifier, using LBFGS with 100 iterations and regularization weight $10^{-3}$. We then use the pre-projection head representations of the test set of the downstream task and evaluate using the aforementioned linear classifer to report the downstream test accuracy.
Using SimCLR for Obtaining Teacher Representation. We conducted an ablation study using a teacher model trained with SimCLR (Chen et al., 2020) instead of Barlow Twins (Zbontar et al., 2021) for CIFAR 10 and CIFAR 100. The experiment steps are similar to the ones in A.1. During the "Training the Teacher Model Using SSL" step , we used the Adam optimizer with batch size 512, learning rate 0.001, and weight decay $10^{-6}$ to train a ResNet18 along with a 2-layer linear projection head for 400 epochs. The projection head included Batch Normalization and ReLU after the first layer, and Batch Normalization after the second layer, projecting to 128 dimensions. Then, we used the pre-projection head representation of the trained model for getting the teacher representation of size 512. The other steps are the same as the one in A.1. Table 6 shows that MKDT consistently outperforms the random subset across all downstream datasets and various sizes of labeled data, highlighting the method’s generalizability to other contrastive learning methods.
@@
Details for KRR-ST Lee et al. (2023) We use the code and hyper-parameters provided in (Lee et al., 2023). As the original paper did not provide the results and the hyperparameters for CIFAR 10, we use the same hyperparameters as CIFAR 100 to distill CIFAR 10. In particular, this invovles using BarlowTwins ResNet-18 as the teacher model as well.
@@ Appendix B Additional Comparison of Random Subset SSL Pre-Training to KRR-ST
This experiment is conducted in the setting of Lee et al. (2023) for pre-training on CIFAR100. This experiment pre-trains on a distilled set / subset of 2% the size of CIFAR100 and then evaluates on downstream tasks by finetuning the entire network with the entire labeled dataset for the downstream task. KRR-ST compares with the random subset baseline by pre-training using SL. However, we use SSL on the random subset as a baseline instead. Here, we show that, in fact, even in the finetuning setting reported by Lee et al. (2023), KRR-ST cannot outperform SSL pre-training on random real images. The baseline reported as ‘Random’ in Lee et al. (2023) refers to SL pre-training as opposed to SSL pre-training.
@@
No pre-training (from Lee et al. (2023)) $64.95$±$\pm 0.21$ $87.34$±$\pm 0.13$ $34.66$±$\pm 0.39$ $19.43$±$\pm 0.14$ $18.46$±$\pm 0.11$ $22.31$±$\pm 0.22$ $58.75$±$\pm 0.41$
Random (Supervised Learning (from Lee et al. (2023)) $65.23$±$\pm 0.12$ $87.55$±$\pm 0.19$ $33.99$±$\pm 0.45$ $19.77$±$\pm 0.21$ $18.18$±$\pm 0.21$ $21.69$±$\pm 0.18$ $59.31$±$\pm 0.27$
KRR-ST (from Lee et al. (2023)) $66.81$±$\pm 0.11$ $88.72$±$\pm 0.11$ $41.54$±$\pm 0.37$ $28.68$±$\pm 0.32$ $25.30$±$\pm 0.37$ $26.39$±$\pm 0.08$ $67.88$±$\pm 0.18$
@@
From Xue et al. (2023), we have that $L_{SSL}$ can be re-written as
@@ Conclusion
It is easy to see, from prior work in hard negative mining for contrastive SSL such as Robinson et al. (2021), that $P_{2}$ will have a different gradient across all batches from $P_{1}$ as $P_{2}$ only contrasts examples within the same class, whereas $P_{1}$ contrasts examples across different classes. Hence, $\theta_{SSL_{1}}\neq\theta_{SSL_{2}}$.
@@
Xue et al. (2023) shows the $L_{SSL}$ can be rewritten as $L_{SSL}=-Tr(2M\theta\theta^{T})+Tr(M^{\prime}\theta^{T}\theta M^{\prime}\theta^{T}\theta)$
@@
From Xue et al. (2023), we also have $\nabla_{\theta}(L_{SSL})=-4\theta M+4\theta M^{\prime}\theta^{T}\theta M^{\prime}$
@@
Appendix E Examples from MKDTand KRR-ST (Lee et al., 2023)
Here, we show examples from CIFAR10, CIFAR100 and TinyImageNet for the data distilled using MKDTand KRR-ST (Lee et al., 2023).
(a) MKDT
(b) KRR-ST
Figure 2: Examples of Distilled Images for CIFAR10
(a) MKDT
(b) KRR-ST
Figure 3: Examples of Distilled Images for CIFAR100
(a) MKDT
(b) KRR-ST
Figure 4: Examples of Distilled Images for TinyImageNet
Appendix F DM and MTT-SSL Experiments
We also compared our methods with two DD methods adapted to SSL. Specifically, we focused on distilling 5% of CIFAR 100. The results all show that those methods do not outperform MKDT.
To adapt DM for SSL, we used the same codebase and hyperparameters as DM (Zhao & Bilen, 2023). However, during distillation, instead of DM’s original approach of sampling random images from a specific class, we modified the process to sample random images from the entire dataset. This adjustment allows us to approximate the behavior of running DM in a label-free setting. The results are shown in 11.
To adapt MTT to SSL, we modified the supervised learning algorithm in getting the trajectory and training the network in the inner loop of the distillation to be the SimCLR SSL algorithm. Due to the high variance of SSL as analyzed in 4.1, distillation fails to converge, and the poor results further shows that MTT with SSL does not work. The results are shown in 12.
Dataset CIFAR100 CIFAR10 TinyImageNet Aircraft CUB2011 Dogs Flowers
CIFAR10 (1% Labels) $7.09\scriptstyle{\pm 0.16}$ $35.98\scriptstyle{\pm 0.98}$ $2.43\scriptstyle{\pm 0.19}$ $2.15\scriptstyle{\pm 0.21}$ $1.19\scriptstyle{\pm 0.13}$ $1.71\scriptstyle{\pm 0.18}$ $1.60\scriptstyle{\pm 0.17}$
CIFAR10 (5% Labels) $14.77\scriptstyle{\pm 0.10}$ $46.41\scriptstyle{\pm 0.35}$ $5.11\scriptstyle{\pm 0.32}$ $4.94\scriptstyle{\pm 0.88}$ $1.57\scriptstyle{\pm 0.31}$ $2.53\scriptstyle{\pm 0.13}$ $3.10\scriptstyle{\pm 0.33}$
CIFAR100 (1% Labels) $8.74\scriptstyle{\pm 0.35}$ $37.75\scriptstyle{\pm 0.87}$ $2.75\scriptstyle{\pm 0.22}$ $2.51\scriptstyle{\pm 0.08}$ $1.20\scriptstyle{\pm 0.11}$ $2.07\scriptstyle{\pm 0.20}$ $2.53\scriptstyle{\pm 0.27}$
CIFAR100 (5% Labels) $17.01\scriptstyle{\pm 0.37}$ $47.32\scriptstyle{\pm 0.38}$ $6.40\scriptstyle{\pm 0.21}$ $5.08\scriptstyle{\pm 0.82}$ $1.99\scriptstyle{\pm 0.13}$ $2.77\scriptstyle{\pm 0.28}$ $4.62\scriptstyle{\pm 0.28}$
Table 11: Results for Distilling with DM with SSL.
Dataset CIFAR100 CIFAR10 TinyImageNet Aircraft CUB2011 Dogs Flowers
CIFAR10 (1% Labels) $4.63\scriptstyle{\pm 0.30}$ $22.06\scriptstyle{\pm 1.86}$ $1.11\scriptstyle{\pm 0.16}$ $1.92\scriptstyle{\pm 0.26}$ $0.77\scriptstyle{\pm 0.11}$ $1.38\scriptstyle{\pm 0.28}$ $1.26\scriptstyle{\pm 0.37}$
CIFAR10 (5% Labels) $7.10\scriptstyle{\pm 0.43}$ $29.04\scriptstyle{\pm 0.58}$ $1.90\scriptstyle{\pm 0.17}$ $2.81\scriptstyle{\pm 0.60}$ $0.88\scriptstyle{\pm 0.08}$ $1.82\scriptstyle{\pm 0.19}$ $1.42\scriptstyle{\pm 0.53}$
CIFAR100 (1% Labels) $4.26\scriptstyle{\pm 0.34}$ $26.06\scriptstyle{\pm 1.58}$ $1.20\scriptstyle{\pm 0.14}$ $1.49\scriptstyle{\pm 0.28}$ $0.87\scriptstyle{\pm 0.26}$ $1.24\scriptstyle{\pm 0.07}$ $1.41\scriptstyle{\pm 0.27}$
CIFAR100 (5% Labels) $9.66\scriptstyle{\pm 0.24}$ $38.16\scriptstyle{\pm 0.57}$ $2.11\scriptstyle{\pm 0.19}$ $2.83\scriptstyle{\pm 0.41}$ $1.06\scriptstyle{\pm 0.04}$ $1.74\scriptstyle{\pm 0.15}$ $2.38\scriptstyle{\pm 0.19}$
Table 12: Results for Distilling with MTT using High Variance SSL Trajectories.
Appendix G Societal Impact